# Interplay between Electroweak Symmetry Breaking and Dark Matter Relic Density

### v2

@@ II.3 Comparing approaches
$\beta$$\equiv$ $A^{I}_{{x^{I}_{f}}\to\infty}-A^{S}_{{x^{S}_{f}}\to\infty}.$ (22)
The sign of $\delta\Omega h^{2}$ is determined by the sign of $\left(\alpha-\beta\right)$. While $\alpha$ quantifies the difference in inverse of yields at freeze-out between the two approaches, $\beta$ is a function of the annihilation cross-sections, with the dependence following from Eqs. (14) and (17). Looking carefully at Eqs. (14) and (17), one may conclude that the quantity $A_{x_{\rm EWSB}\to\infty}$ is the same in both approaches. Further assuming both $\sigma v\right>$’s remain constant over temperature from the decoupling to EWSB, $\beta$ can be further simplified to [^3] In the illustrative model presented below, we have explicitly checked that Eq. (24) correctly predicts the sign of $\delta\Omega h^{2}$ for most (but not all) points in the parameter space. .
$\beta$ $=A^{I}_{{x^{I}_{f}}\to x_{\rm EWSB}}-A^{S}_{{x^{S}_{f}}\to x_{\rm EWSB}}$ (23)
[…]
Quantifying the relative error in this manner highlights the interplay between two significant factors: the annihilation cross-sections and the scalar masses, which together contribute to the shift in the final relic densities between the two approaches.
@@ III.1 Model definition
We extend the SM scalar sector by two real scalar singlets $\phi$ and $\chi$ [17, 30, 33]. The new scalars are stabilized by a $Z_{2}\times Z_{2}^{\prime}$ symmetry such that
@@
where we fix $v_{H}=246$ GeV, and identify $h_{1}$ with the $125$ GeV scalar found at the LHC [34, 35]. Note that the values of these parameters are chosen such that the scalar potential has a minimum both before and after EWSB. All other parameters can be written in terms of these; c.f. App. A.
@@ III.2 Dark sector evolution
[…]
@@ IV.1 Numerical scan
As explained, the upper bound on the mass of $h_{2}$ is chosen to ensure it remains in thermal equilibrium with the SM bath until EWSB occurs. The range considered for the mixing angle $\theta$ is consistent with the latest experimental bounds on heavy Higgs-like scalar mixing [36]. Regarding quartic couplings, we scanned over their typical freeze-out values, ensuring they respect perturbativity and the boundedness from below conditions of App. B. We confirmed that, for values of $\lambda$’s below $10^{-3}$, there are no new features, but occasional numerical instabilities may occur.
To calculate the relic density of $\chi$ in the standard approach, we implemented, as usual, the model described in Sec. III in the electroweak broken phase. We used micrOMEGAs 6.0.5 [32] to numerically integrate the Boltzmann Equation and track the evolution of the yield of $\chi$ from $T=10^{5}$ GeV to $T=10^{-4}$ GeV. The standard relic density was then calculated using Eq. (18). On the other hand, the calculation of the relic density in the improved approach was performed in two steps. In the first step, we implemented the model before EWSB and used micrOMEGAs to integrate the Boltzmann Equation from $T=10^{5}$ GeV to $T=T_{\rm EWSB}$. In the second step, we used the model after EWSB and the previously calculated yield at $T=T_{\rm EWSB}$ as inputs to the darkOmegaNTR() function of micrOMEGAs, allowing us to integrate the Boltzmann Equation for $\chi$ from $T=T_{\rm EWSB}$ to $T=10^{-4}$ GeV. Finally, the improved relic density was then calculated using Eq. (19).
@@
[…]
@@
8.281 467 0.036 $-0.095$ $-0.248$ 5.103 0.118672 0.119259 0.49
[…]
@@ IV.2 Picking a benchmark point
In Fig. 6, we compare the DM yields predicted in the standard and improved approaches, as $\lambda_{H\chi}$ is increased. The top row of plots shows the evolution of the yield of $\chi$ as a function of $1/T$, with $\lambda_{H\chi}$ increasing from left to right. The black solid line represents the yield computed in the improved approach, the red dashed line represents the yield in the standard approach, the gray vertical line indicates the inverse of the EWSB temperature, and the gray horizontal line marks the yield corresponding to the observed relic density. For each value of $\lambda_{H\chi}$, we trace in the bottom row of plots the corresponding temperature evolution of $\langle\sigma v\rangle_{I}$ with a black solid line, and that of $\langle\sigma v\rangle_{S}$ with a red dashed line. It is interesting to note that $\chi$ always decouples from the SM bath earlier than in the improved approach, implying the relation $x_{f}^{S}\lesssim x_{f}^{I}$. Yet, regardless of the value of $\lambda_{H\chi}$, we always observe $\chi$ decoupling from the SM bath when $T>T_{\rm EWSB}$, as expected from the DM mass requirement in Eq. (9).
@@
In the first column of Fig. 6, corresponding to $\lambda_{H\chi}=0.001$, we observe that $\chi$ freezes out with a smaller yield in the improved approach compared to the standard one, leading to an overall negative $\delta\Omega h^{2}$. This behavior can be understood as follows. At freeze-out, the standard and improved yields satisfy $Y_{f}^{S}>Y_{f}^{I}$, which implies a negative value for $\alpha$. On the other hand, before EWSB, we notice that $\langle\sigma v\rangle_{S}\lesssim\langle\sigma v\rangle_{I}$, which implies that $A^{S}_{{x^{I}_{f}}\to\infty}<A^{I}_{{x^{I}_{f}}\to\infty}$ and results in a positive $\beta$. For clarity, the particular values of $\alpha$ and $\beta$ for this benchmark point are marked by the white circular symbol in Fig. 4. A similar conclusion can be drawn for the other two benchmark points highlighted in Fig. 4, which correspond to the $\lambda_{H\chi}$ values used in the middle and right-most columns of Fig. 6. These two points, denoted by a triangle and a diamond in Fig. 4, exhibit positive $\alpha$. This arises from the fact that, before EWSB, the total thermal averaged cross-sections follow $\langle\sigma v\rangle_{S}>\langle\sigma v\rangle_{I}$, leading to a faster depletion of DM particles in the standard approach, and resulting in $Y_{f}^{S}<Y_{f}^{I}$. Due to the aforementioned tension between relevant parameters, $\beta$ remains negative in both cases. As a result, $\delta\Omega h^{2}$ turns out to be positive, in agreement with the values of $\delta\Omega h^{2}$ quoted on each row, and with the position of the solid black line (improved yield) lying above the red dashed one (standard yield). Note that for $\lambda_{H\chi}=4.9$, the improved relic density matches the observed value $\Omega h^{2}|_{\rm Obs}=0.12$ at the $3\sigma$ level, whereas the relic density computed in the standard approach does not. In fact, the standard relic density deviates from $\Omega h^{2}|_{\rm Obs}$ by $42\sigma$.
@@ V Conclusions
It is important to note that these heavy DM particles are not excluded by experiment. The ongoing and planned direct and indirect detection experiments (see Ref. [1] for a comprehensive review) will most certainly give rise to an increase (or even exclusion) in the upper bound of the DM mass in this simple scenario with two singlets and two independent symmetries. However, it is easy to extend the model with, for example, an extra singlet, such that the couplings responsible for direct and indirect detection are not the same as the ones that determine the value of the relic density. We defer a detailed investigation of such extensions to future work.
The study also builds a strong case for testing TeV-scale phenomena in upcoming high-energy particle colliders such as the FCC-hh [40]. In fact, although models for which our approach is important have very heavy DM candidates, they could in principle be probed at future colliders. In this case the approach could prove instrumental in the reconstruction of a given potential from an extended scalar sector
@@ Appendix B Boundedness from below
For the scalar potential studied in this paper to be bounded from below, we must ensure that [41, 42]:
@@ Appendix C Direct Detection
The DM-nucleon spin-independent elastic scattering cross-section ($\sigma_{\mathrm{SI}}$), in this model reads [43]
@@
where $f_{N}\approx 0.308$ is the Higgs-nucleon effective coupling [44], $m_{N}\approx 0.938$ GeV is the nucleon mass, and

### v3

@@ II.3 Comparing approaches
$\beta$ $\equiv$ $A^{I}_{{y^{I}_{f}}\to\infty}-A^{S}_{{y^{S}_{f}}\to\infty}.$ (23)
The sign of $\delta\Omega h^{2}$ is determined by the sign of $\left(\alpha-\beta\right)$. While $\alpha$ quantifies the difference in inverse of yields at freeze-out between the two approaches, $\beta$ is a function of the annihilation cross-sections, with the dependence following from Eqs. (15) and (18). Looking carefully at Eqs. (15) and (18), one may conclude that the quantity $A_{y_{\rm EWSB}\to\infty}$ is the same in both approaches. Quantifying the relative error in this manner highlights the interplay between two significant factors: the annihilation cross-sections and the scalar masses, which together contribute to the shift in the final relic densities between the two approaches.
@@ III.1 Model definition
We extend the SM scalar sector by two real scalar singlets $\phi$ and $\chi$ Cline:2013gha ; Steigman:2012nb ; Basak:2021tnj . The new scalars are stabilized by a $Z_{2}\times Z_{2}^{\prime}$ symmetry such that
@@
where we fix $v_{H}=246$ GeV, and identify $h_{1}$ with the $125$ GeV scalar found at the LHC ATLAS:2012yve ; CMS:2012qbp . Note that the values of these parameters are chosen such that the scalar potential has a minimum both before and after EWSB. All other parameters can be written in terms of these; c.f. App. A.
@@ III.2 Dark sector evolution
If both components were to freeze out after EWSB, then, the first line of Eq. (18) would vanish, leading to $A^{I}_{{y_{1}}\to y_{2}}=A^{S}_{{y_{1}}\to y_{2}}$ of Eq. (15), and the two approaches would coincide. This would only be achieved if the mass hierarchies were $m_{\chi},m_{\phi}\lesssim\ 4\,{\rm TeV}$, due to the reason mentioned in Eq. (10). To simplify the analysis conservatively, in our study, we impose $m_{\phi}\lesssim\,1\ {\rm TeV}$, so that $\phi$ remains in the thermal bath before EWSB. As a result, $Y_{\phi}$ always traces $\overline{Y}_{\phi}$ until EWSB occurs. On the other hand, we require $\widetilde{m}_{\chi}\gtrsim 4\ {\rm TeV}$, such that $Y_{\chi}$ saturates before EWSB. Thus, the treatment of $\chi$ in the standard and improved approaches will have a noticeable physical consequence.
@@ IV.1 Numerical scan
As explained, the upper bound on the mass of $h_{2}$ is chosen to ensure it remains in thermal equilibrium with the SM bath until EWSB occurs. The range considered for the mixing angle $\theta$ is consistent with the latest experimental bounds on heavy Higgs-like scalar mixing ATLAS:2021pkb . Regarding quartic couplings, we scanned over their typical freeze-out values, ensuring they respect perturbativity and the boundedness from below conditions of App. B. We confirmed that, for values of $\lambda$’s below $10^{-3}$, there are no new features, but occasional numerical instabilities may occur.
To calculate the relic density of $\chi$ in the standard approach, we implemented, as usual, the model described in Sec. III in the electroweak broken phase. We used micrOMEGAs 6.0.5 Alguero:2023zol to numerically integrate the Boltzmann Equation and track the evolution of the yield of $\chi$ from $T=10^{5}$ GeV to $T=10^{-4}$ GeV. The standard relic density was then calculated using Eq. (19). On the other hand, the calculation of the relic density in the improved approach was performed in two steps. In the first step, we implemented the model before EWSB and used micrOMEGAs to integrate the Boltzmann Equation from $T=10^{5}$ GeV to $T=T_{\rm EWSB}$. In the second step, we used the model after EWSB and the previously calculated yield at $T=T_{\rm EWSB}$ as inputs to the darkOmegaNTR() function of micrOMEGAs, allowing us to integrate the Boltzmann Equation for $\chi$ from $T=T_{\rm EWSB}$ to $T=10^{-4}$ GeV. Finally, the improved relic density was then calculated using Eq. (20).
@@
In Fig. 3, we scan over the parameter space and plot, in gray, the relative difference ($\delta\Omega h^{2}$) between the standard and improved approaches as a function of the dark matter mass, over the range 11001–100 TeV. Here, the gray vertical line marks the inverse of the EWSB temperature. Points with relic density computed in the improved approach and lying within the $3\sigma$ range of the PLANCK measurement Planck:2018vyg are shown in red. As expected, $\delta\Omega h^{2}$ begins to deviate noticeably from zero as $m_{\chi}$ approaches 4 TeV — the approximate lower bound on the dark matter mass required for it to decouple from the SM bath before electroweak symmetry breaking, as discussed in Sec. II. A subset of the red points, marked with green, satisfy the additional requirement that the direct detection (DD) cross-section lies below the current experimental limits Bhoonah:2020fys ; LZ:2024psa . We provide an explicit formula for the spin-independent scattering cross-section in App. C. For $m_{\chi}\lesssim 10\ {\rm TeV}$, we applied the latest DD exclusion limits on the spin-independent scattering cross-section off nuclei from the 2024 LUX-ZEPLIN (LZ) experiment LZ:2024zvo . We found that the green points tend to favour small values of $\delta\Omega h^{2}$, reflecting the fact that in this simplified model, both the relic density and the DD cross-section are controlled by the scalar quartic couplings $\lambda_{H\chi}$ or $\lambda_{\phi\chi}$. These couplings need to be large to achieve the observed relic density, but small enough to pass DD constraints. However, we expect these exclusions to be less restrictive in more complete models, where the relic abundance and DD channels are not necessarily correlated. In Table. 1, we show two red benchmark points with $m_{\chi}\lesssim\ 10\ {\rm TeV}$ with non-zero $\delta\Omega h^{2}$ that have DD cross-sections below the 2024 LZ bounds LZ:2024zvo .
@@
8.281 467 0.036 0.095-0.095 0.248-0.248 5.103 0.118672 0.119259 0.49
Table 1: Benchmark points with $m_{\chi}\lesssim\,10\ {\rm TeV}$ with finite $\delta\Omega h^{2}$ and direct detection cross-sections below LZ limits.
@@ IV.2 Picking a benchmark point
In Fig. 6, we compare the DM yields predicted in the standard and improved approaches, as $\lambda_{H\chi}$ is increased. The top row of plots shows the evolution of the yield of $\chi$ as a function of $1/T$, with $\lambda_{H\chi}$ increasing from left to right. The black solid line represents the yield computed in the improved approach, the red dashed line represents the yield in the standard approach, the gray vertical line indicates the inverse of the EWSB temperature, and the gray horizontal line marks the yield corresponding to the observed relic density. For each value of $\lambda_{H\chi}$, we trace in the bottom row of plots the corresponding temperature evolution of $\langle\sigma v\rangle_{I}$ with a black solid line, and that of $\langle\sigma v\rangle_{S}$ with a red dashed line. It is interesting to note that $\chi$ always decouples from the SM bath earlier than in the improved approach, implying the relation $y_{f}^{S}\lesssim y_{f}^{I}$. Yet, regardless of the value of $\lambda_{H\chi}$, we always observe $\chi$ decoupling from the SM bath when $T>T_{\rm EWSB}$, as expected from the DM mass requirement in Eq. (10).
@@
In the first column of Fig. 6, corresponding to $\lambda_{H\chi}=0.001$, we observe that $\chi$ freezes out with a smaller yield in the improved approach compared to the standard one, leading to an overall negative $\delta\Omega h^{2}$. This behavior can be understood as follows. At freeze-out, the standard and improved yields satisfy $Y_{f}^{S}>Y_{f}^{I}$, which implies a negative value for $\alpha$. On the other hand, before EWSB, we notice that $\langle\sigma v\rangle_{S}\lesssim\langle\sigma v\rangle_{I}$, which implies that $A^{S}_{{y^{I}_{f}}\to\infty}<A^{I}_{{y^{I}_{f}}\to\infty}$ and results in a positive $\beta$. For clarity, the particular values of $\alpha$ and $\beta$ for this benchmark point are marked by the white circular symbol in Fig. 4. A similar conclusion can be drawn for the other two benchmark points highlighted in Fig. 4, which correspond to the $\lambda_{H\chi}$ values used in the middle and right-most columns of Fig. 6. These two points, denoted by a triangle and a diamond in Fig. 4, exhibit positive $\alpha$. This arises from the fact that, before EWSB, the total thermal averaged cross-sections follow $\langle\sigma v\rangle_{S}>\langle\sigma v\rangle_{I}$, leading to a faster depletion of DM particles in the standard approach, and resulting in $Y_{f}^{S}<Y_{f}^{I}$. Due to the aforementioned tension between relevant parameters, $\beta$ remains negative in both cases. As a result, $\delta\Omega h^{2}$ turns out to be positive, in agreement with the values of $\delta\Omega h^{2}$ quoted on each row, and with the position of the solid black line (improved yield) lying above the red dashed one (standard yield). Note that for $\lambda_{H\chi}=4.9$, the improved relic density matches the observed value $\Omega h^{2}|_{\rm Obs}=0.12$ at the $3\sigma$ level, whereas the relic density computed in the standard approach does not. In fact, the standard relic density deviates from $\Omega h^{2}|_{\rm Obs}$ by $42\sigma$.
@@ V Conclusions
It is important to note that these heavy DM particles are not excluded by experiment. The ongoing and planned direct and indirect detection experiments (see Ref. Cirelli:2024ssz for a comprehensive review) will most certainly give rise to an increase (or even exclusion) in the upper bound of the DM mass in this simple scenario with two singlets and two independent symmetries. However, it is easy to extend the model with, for example, an extra singlet, such that the couplings responsible for direct and indirect detection are not the same as the ones that determine the value of the relic density. We defer a detailed investigation of such extensions to future work.
The study also builds a strong case for testing TeV-scale phenomena in upcoming high-energy particle colliders such as the FCC-hh Benedikt:2022kan . In fact, although models for which our approach is important have very heavy DM candidates, they could in principle be probed at future colliders. In this case the approach could prove instrumental in the reconstruction of a given potential from an extended scalar sector
@@ Appendix B Boundedness from below
For the scalar potential studied in this paper to be bounded from below, we must ensure that Kannike:2012pe ; Habibolahi:2022rcd :
@@ Appendix C Direct Detection
The DM-nucleon spin-independent elastic scattering cross-section ($\sigma_{\mathrm{SI}}$), in this model reads Cerdeno:2010jj
@@
where $f_{N}\approx 0.308$ is the Higgs-nucleon effective coupling Hoferichter:2017olk , $m_{N}\approx 0.938$ GeV is the nucleon mass, and